69Cu decays by emitting a β- particle. This decay process has a half-life of 2.80 min. Answer the following questions about this process and report all answers to three significant figures.

1. Determine the initial activity (in dis/min, do not input units) of a 8.68 mg sample of this isotope. Report your answer to three significant figures in scientific notation.

2. Determine the time (in min) that it takes for the activity of 69Cu to decrease to 3.00×106 dis/min.
Answer #1

Solution-
Sol.

(A) As we know that the half life = t1/2 = 2.80 min

So , Decay constant = K

= 0.693 / t1/2 = 0.693 / 2.80 = 0.2475 min-1

So the , Mass of Cu = 8.68 mg = 8.68 * 10^-3 g

And the molar Mass of Cu = 69 g/mol

Now the, initial number of nuclei present

= ( Mass of Cu / Molar Mass of Cu ) × Avogadro's no.

= ( ( 8.68 * 10^-3 ) / 69 ) * 6.022 * 10^23

= 7.57 * 10^19

Hence the , initial activity of Cu = A°

= initial number of nuclei present * Decay constant

= (7.57 *10^19)* 0.2475

= 1.87 * 10^19 dis/min

(B)
A = activity of Cu after time t =3.00*10^6 dis/min

Therefore , ln ( A° / A ) = K t

ln((1.87 * 10^19 )/(3.00*10^6)) = 0.2475 t

0.2475 t = 29.46

t = 119.03 min